OpenStudy (anonymous):

Write an equation of the parabola with vertex (0, 3) and focus (-4, 3).

OpenStudy (anonymous):

\[(x+4)^2+(y-3)^2=(x-4)^2\] I think

OpenStudy (amistre64):



OpenStudy (amistre64):

y^2 = 4ax -y^2 = 4(4)x -(1/16) y^2 = x -(1/16) (y-3)^2 = x

OpenStudy (amistre64):

im right :) http://www.wolframalpha.com/input/?i=-%28y-3%29^2+%3D+16x

OpenStudy (anonymous):

Thanks =)

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

lol .... thats cheating

OpenStudy (anonymous):

Its the same equation, one is just solved for x lol